# Less but Better: Enabling Generalized Zero-shot Learning Towards Unseen Domains by Intrinsic Learning from Redundant LLM Semantics

### v4

@@ III-D Adversarial construction of common feature space
Here, $x_{i}$ denotes the $i-th$ training sample, $k$ the total number of training samples, and $y_{i}^{d}$ and $y_{i}^{c}$ represent the domain and class labels of $x_{i}$ respectively. $1[\cdot]$ serves as the indicator function.
@@
III-E Meta semantic refinement in common space
@@
III-E1 Inter-class similarity alignment (ISA)
Due to the presence of non-intrinsic semantics, the inter-class similarity obtained directly from the original semantics differs from the inter-class similarity obtained in the data-driven feature space $\Psi$. Therefore, if we can refine the semantics based on the generalizable inter-class similarity in the common space $\Psi$, we can indirectly remove the non-intrinsic semantics. Specifically, we construct the inter-class similarity leveraging the dark knowledge contained in the output of classifier $C_{s}$. The reason for using dark knowledge is that studies have shown that irrelevant information, like background details, still persists in the feature space [26]. In contrast, dark knowledge, found in the classifier’s final layer, could better represent the essential information in the feature space related to class semantics because it bases the perspective of the final classification. Therefore, we can leverage the class similarity relationships calculated by dark knowledge and the similarity relationships calculated by semantics to establish the refinement constraint. The content of this part is displayed in the lower left corner of Fig. 6.
@@ III-E1 Inter-class similarity alignment (ISA)
$cos$ represents the cosine similarity, and $M$ represents the similarity of each class dark knowledge prototype in $\Psi$.
Next, we come to the semantic similarity part. We build a semantic encoder $E_{s}$ for semantic refinement. It should be noted that our approach is applicable to different kinds of semantics. If the semantics are text extracted from LLMs, the initial $E_{s}$ could be a pre-trained sentence-Bert [27] that maps a paragraph to a fixed-length vector. If the semantics are attributes provided by human experts, $E_{s}$ could be a multi-layer perceptron network. With the semantic encoder $E_{s}$, the refined semantics can be expressed as:
@@
After the non-intrinsic semantics are eliminated, the inter-class similarity of semantics will become closer to the similarity in the feature space, because the common feature space does not include non-intrinsic semantic information. Therefore, if $M$ and $N$ are closely aligned, the semantic redundancy could be mitigated. The related loss function could be expressed as:
@@
III-E2 Unseen-class meta generation (UMG)
@@
In order to simulate the ZSL generation process, we implement our semantic refinement within the meta-learning framework. The content of this part is displayed in the lower right corner of Fig. 6. Differing from traditional machine learning, the fundamental unit of meta-learning is a task, not a sample. For each task $\mathcal{T}_{i}=\{\mathcal{T}_{su},\mathcal{T}_{qu}\}$, drawn from the task distribution $p(\mathcal{T})$, the support set $\mathcal{T}_{su}$ comprises $N_{su}$ seen classes, while the query set $\mathcal{T}_{qu}$ includes $N_{qu}$ seen classes. To mirror the ZSL process, the classes within the support set and the query set are disjoint. Generators that fit the features of support classes with the semantics refined by $E_{s}$ should be capable of generating reliable features of query classes. For each pair of support and query sets, we can generate the refined support set semantics using $E_{s}$, and then fit the support set and validate it with the query set for these respective tasks. Therefore, we could employ the loss to optimize semantic encoder $E_{s}$ to ensure it preserve intrinsic semantics beneficial to generation.
Specifically, within each task, we utilize $E_{s}$ to obtain refined support set semantics $A^{\circ}_{su}$ and then train a generator from support classes. After that, fake query features are synthesized by refined query semantics and the above generator. Since real query features are at hand, loss of the query set could be calculated. We employ the loss of each task’s query set to evaluate the effectiveness of $E_{s}$. If the loss on each task is relatively small, this indicates that the current $E_{s}$ is near optimality. Otherwise, we could optimize $E_{s}$ using the gradient of this loss with respect to $E_{s}$. In order to make the loss of the query set have an effective gradient with respect to $E_{s}$, a linear least squares generator (LLSG) $z$ is leveraged here, where the independent variable is refined class semantics and the dependent variable is the prototype of class features, as only a linear generator can obtain the following analytic solution:
@@
The fake query feature prototypes $\tilde{U}_{qu}$ produced by $z^{*}$ could be presented as:
@@
It’s easy to find (13) is only differential to the parameter of $E_{s}$, and the meta-loss can be obtained from each task and optimize $E_{s}$.
@@
Considering the similarity constraint, the final loss of meta semantic refinement could be represented as:
@@
With the optimized $E_{s}$, we could obtain the refined semantics $A^{\circ}$, including the seen and unseen classes, which are $A^{\circ}_{s}=E_{s}(A_{s})$ and $A^{\circ}_{u}=E_{s}(A_{u})$.
III-E3 Feature generation with refined semantics
In this stage, we would like to train a formal generator that can synthesize features in the common space under the condition of $A^{\circ}$. Then, the fake unseen features and real seen features are leveraged together to obtain the final classifier. Specifically, we utilized WGAN-GP [28] as our formal generator and the training process is shown as follows:
@@ III-E3 Feature generation with refined semantics
where $E$ represents the empirical expectation, $G$ and $D$ represent the generator and discriminator of WGAN-GP, $f$ represents the real features, and $\tilde{f}$ represents the generated features, $\beta$ is a hyperparameter.
The generator is trained on seen class data conditioned on the refined semantics $A^{\circ}$. Because of the refined semantics, the information asymmetry between semantic space and feature space is solved beforehand, and this decreases the difficulty of the mapping of the generator.
After training the generator, we could obtain the fake unseen features through:
@@
where $\epsilon$ is the random sampled Gaussian noise.
@@
III-F The CDGZSL inference procedure
In the last subsection, we would like to introduce the inference process of MDASR. For testing data $x_{t}$, which comes from the unseen domain and could belong to seen and unseen classes, we first map it to the common feature space with data encoder $E_{d}$, and then predict its label with CDGZSL classifier $C_{c}$. The whole process could be formalized as follows:
@@ III-F The CDGZSL inference procedure
Because $C_{c}$ is trained by both real seen features and fake unseen features with less effection of semantic redundancy, the domain shift problem could be further alleviated, and we could gain a more accurate classifier of CDGZSL.
IV Experiment
@@
In this section, we have verified the proposed MDASR on two public datasets, Office-Home[29] and Mini-DomainNet[30]. At the same time, we have also carried out a series of comparative experiments, and ablation studies to further demonstrate the effectiveness of the proposed method.
IV-A Datasets and relative semantics
@@
Different from the existing CDZSL studies which utilize word vector[31] to serve as semantics, we employ LLMs to extract semantics for the two datasets. In this way, we could validate the performance of our proposed methodology in refining LLM semantics. Due to the relatively small number of categories in Office-Home, in addition to LLM semantics, we also provide a manually annotated class-attribute matrix, which is the traditional form of semantics for ZSL, used for comparison with the semantics provided by LLM. For the convenience of other researchers, we provide all the above semantics on the Internet. [^1] https://github.com/chunhuiz/Semantics-of-OfficeHome-and-MiniDomainNet
IV-A1 Office-Home
@@
The Office-Home dataset is a popular dataset used for multi-domain transfer learning tasks. It contains about 15,500 images that belong to 65 different object classes. The images are sourced from four significantly different domains: Art, Clipart, Product, and Real-world. As mentioned earlier, in the semantic annotations of the Office-Home dataset, we employed two forms: LLM annotation and manual annotation. For LLM annotation, as mentioned in Section III.B, we employed GPT-3.5 [32] as the LLM to obtain descriptive paragraphs as semantics for each class. For manual annotation, we first predefined 21 attributes, including the material, size, color, etc., and annotated the 65 classes, forming a $65\times 21$ class-attribute matrix, with elements being 0 or 1.
IV-A2 Mini-DomainNet
@@
The Mini-DomainNet is a large-scale dataset which simplified from the DomainNet [33] dataset. It contains about 140000 images that belong to 126 different object classes. The images are sourced from four diverse domains: Art, Clipart, Sketch, and Real. Compared to Office-Home, Mini-DomainNet has more classes, consumes more computational resources, and poses greater challenges. Due to the large number of classes in this dataset, we did not have the resources to perform manual annotations, so we only employed the LLM for semantic annotation.
TABLE II: The CDGZSL results on Office-Home dataset
{tblr}
colspec=c, cell11 = r=2, cell12 = c=2,r=2, cell14 = c=3, cell17 = c=3, cell110 = c=3, cell31 = r=8, cell32 = r=3, cell62 = c=2, cell72 = c=2, cell82 = r=2, cell111 = r=8, cell112 = r=3, cell142 = c=2, cell152 = c=2, cell162 = r=2, cell182 = c=2, vline2 = 1-18, hline1,3,11,19 = -, hline2 = 4-12, hline4-5,9,13,19-20 = 3, hline6-8,10,11-12,14-15,16-18,21-22 = 2-3, Semantics & Methods Art Clipart Product
S U H S U H S U H
LLM text CuMix img-only 59.43 0 0 50.91 3.28 6.16 66.21 0.72 1.42
two-level 57.37 1.85 3.58 50.77 3.61 6.74 66.88 0.70 1.38
original 57.83 2.46 4.71 51.19 3.94 7.31 65.09 1.80 3.50
S-AGG 55.09 1.23 2.40 49.90 7.56 13.13 63.38 2.52 4.84
SEIC 56.64 2.48 4.75 48.49 8.55 14.53 63.19 3.97 7.47
f-CLSWGAN original 56.90 43.15 49.08 53.56 28.26 36.99 72.94 34.46 46.80
PCA 57.50 38.88 46.39 53.53 25.98 34.98 73.61 27.07 39.58
MDASR 55.10 50.62 52.76 52.67 36.18 42.89 73.19 38.26 50.25
Attribute CuMix img-only 43.81 4.93 8.86 41.10 2.96 5.52 56.49 2.52 4.82
two-level 43.45 5.56 9.85 39.35 4.60 8.23 55.17 6.49 11.61
original 44.28 6.17 10.83 38.95 7.23 12.19 53.16 9.38 15.94
S-AGG 45.31 5.60 9.96 37.23 8.89 14.35 50.41 7.94 13.71
SEIC 43.50 6.27 10.96 41.41 5.26 9.33 54.70 6.85 12.17
f-CLSWGAN original 54.85 35.18 42.87 53.42 28.94 37.54 72.91 32.12 44.59
PCA 57.85 31.48 40.77 53.92 29.29 37.95 72.94 22.02 33.82
MDASR 59.10 38.27 46.45 53.28 30.26 38.60 72.99 40.07 51.74
@@
IV-B Experimental details
In CDGZSL, there are both seen and unseen domains, as well as seen and unseen classes. For Office-Home, the allocation of seen and unseen domains is as follows: Art, Clipart, and Product are each designated as the unseen domain in turn. When a domain is designated as the unseen domain, the other three domains serve as the seen domains. Because we employ Resnet-50 pretrained on Imagenet[34] as the initial parameters for the data encoder $E_{d}$, only the Real-world domain is not used as the unseen domain (since it is more similar to Imagenet) among these four domains. Additionally, the division of seen and unseen classes is as follows: we select 5 classes that do not appear in Imagenet as unseen classes, 5 classes as validation, and the remaining 55 classes as seen classes. For Mini-DomainNet, similarly, aside from the Real domain, the other three domains each take a turn as the unseen domain once. Also, we select 13 classes that do not appear in Imagenet as unseen classes, 13 classes as validation, and the remaining classes as seen classes.
For the various networks used in the experiments, the specific details are as follows: The data encoder $E_{d}$ is the encoder of Resnet50, with initial parameters trained on Imagenet; The domain classifier $C_{d}$, seen class classifier $C_{s}$, and the final CDGZSL classifier $C_{c}$ are all single-layer fully connected layers followed by Softmax. The semantic encoder $E_{s}$ is divided into two cases: When the semantics are text generated by LLM, $E_{s}$ is sentence-Bert [27], with initial parameters using all-MiniLM-L6-v2, and only the last two layers are fine-tuned during training. When the semantics are manually annotated attributes, $E_{s}$ is a two-layer fully connected layer with an output dimension of 15. The generator and discriminator of WGAN-GP are also both two-layer fully connected layers. The activation function of all the above-mentioned fully connected layers is ReLU. For the optimizers and various hyperparameters used in training, the specific details are as follows: $\lambda=1$, $T=5$, $\alpha=0.1$, $\beta=1$, the optimizer is Adam, and the learning rate is $2\times 10^{-4}$.

### v5

@@ 3.4 Adversarial construction of common feature space
Here, $x_{i}$ denotes the $i-th$ training sample, $k$ the total number of training samples, and $y_{i}^{d}$ and $y_{i}^{c}$ the domain and class labels of $x_{i}$ respectively. $1[\cdot]$ serves as the indicator function.
@@
3.5 Meta semantic refinement in common space
@@
3.5.1 Inter-class similarity alignment (ISA)
Due to the presence of non-intrinsic semantics, the inter-class similarity obtained directly from the original semantics differs from the inter-class similarity obtained in the data-driven feature space $\Psi$. Therefore, if we can refine the semantics based on the generalizable inter-class similarity in the common space $\Psi$, we can indirectly remove the non-intrinsic semantics. Specifically, we construct the inter-class similarity leveraging the dark knowledge contained in the output of classifier $C_{s}$. The reason for using dark knowledge is that studies have shown that irrelevant information, like background details, still persists in the feature space (Cheng et al., 2020). In contrast, dark knowledge, found in the classifier’s final layer, could better represent the essential information in the feature space related to class semantics because it bases the perspective of the final classification. Therefore, we can leverage the class similarity relationships calculated by dark knowledge and the similarity relationships calculated by semantics to establish the refinement constraint. The content of this part is displayed in the lower left corner of Fig. 6.
@@ 3.5.1 Inter-class similarity alignment (ISA)
$cos$ represents the cosine similarity, and $M$ represents the similarity of each class of dark knowledge prototype in $\Psi$.
Next, we come to the semantic similarity part. We build a semantic encoder $E_{s}$ for semantic refinement. It should be noted that our approach is applicable to different kinds of semantics. If the semantics are text extracted from LLMs, the initial $E_{s}$ could be a pre-trained sentence-Bert (Reimers and Gurevych, 2019) that maps a paragraph to a fixed-length vector. If the semantics are attributes provided by human experts, $E_{s}$ could be a multi-layer perceptron network. With the semantic encoder $E_{s}$, the refined semantics can be expressed as:
@@
After the non-intrinsic semantics are eliminated, the inter-class similarity of semantics will become closer to the similarity in the feature space, because the common feature space does not include non-intrinsic semantic information. Therefore, if $M$ and $N$ are closely aligned, the semantic redundancy could be mitigated. The related loss function can be expressed as:
@@
3.5.2 Unseen-class meta generation (UMG)
@@
In order to simulate the ZSL generation process, we implement our semantic refinement with meta-learning. The content of this part is displayed in the lower right corner of Fig. 6. Differing from traditional machine learning, the fundamental unit of meta-learning is a task, not a sample. For each task $\mathcal{T}_{i}=\{\mathcal{T}_{su},\mathcal{T}_{qu}\}$, drawn from the task distribution $p(\mathcal{T})$, the support set $\mathcal{T}_{su}$ comprises $N_{su}$ seen classes, while the query set $\mathcal{T}_{qu}$ includes $N_{qu}$ seen classes. To mirror the ZSL process, the classes within the support set and the query set are disjoint. Generators that fit the features of support classes with the semantics refined by $E_{s}$ should be capable of generating reliable features of query classes. For each pair of support and query sets, we can generate the refined support set semantics using $E_{s}$, and then fit the support set and validate it with the query set for these respective tasks. Therefore, we could employ the loss to optimize semantic encoder $E_{s}$ to ensure it preserve intrinsic semantics beneficial to generation.
Different with exsiting meta-learning work, within each task, we utilize $E_{s}$ to obtain refined support set semantics $A^{\circ}_{su}$ and then train a generator from support classes. After that, fake query features are synthesized by refined query semantics and the above generator. Since real query features are at hand, loss of the query set can be calculated. We employ the loss of each task’s query set to evaluate the effectiveness of $E_{s}$. If the loss on each task is relatively small, this indicates that the current $E_{s}$ is near the optimality. Otherwise, we could optimize $E_{s}$ using the gradient of this loss with respect to $E_{s}$. In order to make the loss of the query set have an effective gradient with respect to $E_{s}$, a linear least squares generator (LLSG) $z$ is leveraged here, where the independent variable is the refined class semantics and the dependent variable is the prototype of class features, as only a linear generator can obtain the following analytic solution:
@@
The fake query feature prototypes $\tilde{U}_{qu}$ produced by $z^{*}$ can be presented as:
@@
It is easy to find (13) is only differential to the parameter of $E_{s}$, and the meta-loss can be obtained from each task and then optimizing $E_{s}$.
@@
Considering the similarity constraint, the final loss of meta semantic refinement can be represented as:
@@
With the optimized $E_{s}$, we can obtain the refined semantics $A^{\circ}$, including the seen and unseen classes, which are $A^{\circ}_{s}=E_{s}(A_{s})$ and $A^{\circ}_{u}=E_{s}(A_{u})$.
3.5.3 Feature generation with refined semantics
At this stage, we will train a formal generator that can synthesize features in the common space under the condition $A^{\circ}$. Then, the fake unseen features and real seen features are leveraged together to obtain the final classifier. Specifically, we utilized WGAN-GP (Gulrajani et al., 2017) as our formal generator and the training process is shown as follows:
@@ 3.5.3 Feature generation with refined semantics
where $E$ represents the empirical expectation, $G$ and $D$ represent the generator and discriminator of WGAN-GP, $f$ represents the real features, and $\tilde{f}$ represents the generated features, and $\beta$ is a hyperparameter.
The generator is trained on seen class data conditioned on the refined semantics $A^{\circ}$. Because of the refined semantics, the information asymmetry between semantic space and feature space is solved beforehand, and this decreases the difficulty of the generator mapping.
After training the generator, we obtain the fake unseen features through:
@@
where $\epsilon$ is the randomly sampled Gaussian noise.
@@
3.6 The CDGZSL inference procedure
In this last subsection, we will introduce the inference process of MDASR. For testing data $x_{t}$, which comes from the unseen domain and could belong to seen and unseen classes, we first map it to the common feature space with data encoder $E_{d}$, and then predict its label with CDGZSL classifier $C_{c}$. The whole process can be formalized as follows:
@@ 3.6 The CDGZSL inference procedure
Because $C_{c}$ is trained by both real seen features and fake unseen features with less semantic redundancy, the domain shift problem could be further alleviated, and we could gain a more accurate classifier of CDGZSL.
4 Experiment
@@
In this section, we have verified the proposed MDASR on two public datasets, Office-Home(Venkateswara et al., 2017) and Mini-DomainNet(Zhou et al., 2021). At the same time, we have also carried out a series of comparative experiments, and ablation studies to further demonstrate the effectiveness of the proposed method.
4.1 Datasets and relative semantics
@@
Different from the existing CDZSL studies which utilize word vector(Church, 2017) to serve as semantics, we employ LLMs to extract semantics for the two datasets. In this way, we can validate the performance of our proposed methodology in refining LLM semantics. Due to the relatively small number of categories in Office-Home, in addition to LLM semantics, we also provide a manually annotated class-attribute matrix, which is the traditional form of semantics for ZSL, used for comparison with the semantics provided by LLM. For the convenience of other researchers, we provide all the above semantics to the public. [^1] https://github.com/chunhuiz/Semantics-of-OfficeHome-and-MiniDomainNet
4.1.1 Office-Home
@@
The Office-Home dataset is a popular dataset used for multi-domain transfer learning tasks. It contains about 15,500 images that belong to 65 different object classes. The images are sourced from four significantly different domains: Art, Clipart, Product, and Real-world. As mentioned earlier, in the semantic annotations of the Office-Home dataset, we employed two forms: LLM annotation and manual annotation. For LLM annotation, as mentioned in Section III.B, we employed GPT-3.5 (Kocoń et al., 2023) as the LLM to obtain descriptive paragraphs as semantics for each class. For manual annotation, we first predefined 21 attributes, including the material, size, color, etc., and annotated the 65 classes, forming a $65\times 21$ class-attribute matrix, with elements being 0 or 1.
4.1.2 Mini-DomainNet
@@
The Mini-DomainNet is a large-scale dataset which is simplified from the DomainNet (Peng et al., 2019) dataset. It contains about 140000 images that belong to 126 different object classes. The images are sourced from four diverse domains: Art, Clipart, Sketch, and Real. Compared to Office-Home, Mini-DomainNet has more classes, consumes more computational resources, and poses greater challenges. Due to the large number of classes in this dataset, we did not perform manual annotations, so we only employed the LLM for semantic annotation.
Table 3: The CDGZSL results on Office-Home dataset
Semantics Methods Art Clipart Product S U H S U H S U H LLM text CuMix img-only 59.43 0 0 50.91 3.28 6.16 66.21 0.72 1.42 two-level 57.37 1.85 3.58 50.77 3.61 6.74 66.88 0.70 1.38 original 57.83 2.46 4.71 51.19 3.94 7.31 65.09 1.80 3.50 S-AGG 55.09 1.23 2.40 49.90 7.56 13.13 63.38 2.52 4.84 SEIC 56.64 2.48 4.75 48.49 8.55 14.53 63.19 3.97 7.47 ZeroDiff 45.20 27.16 33.93 37.01 13.16 19.41 63.73 19.86 30.28 VADS 43.85 16.05 23.50 40.11 25.01 30.80 60.51 2.89 5.51 MDASR 55.10 50.62 52.76 52.67 36.18 42.89 73.19 38.26 50.25 Attribute CuMix img-only 43.81 4.93 8.86 41.10 2.96 5.52 56.49 2.52 4.82 two-level 43.45 5.56 9.85 39.35 4.60 8.23 55.17 6.49 11.61 original 44.28 6.17 10.83 38.95 7.23 12.19 53.16 9.38 15.94 S-AGG 45.31 5.60 9.96 37.23 8.89 14.35 50.41 7.94 13.71 SEIC 43.50 6.27 10.96 41.41 5.26 9.33 54.70 6.85 12.17 ZeroDiff 45.55 17.90 25.70 38.61 24.34 29.86 56.93 36.10 44.19 VADS 35.25 27.78 31.07 40.67 18.09 25.04 61.62 14.08 22.92 MDASR 59.10 38.27 46.45 53.28 30.26 38.60 72.99 40.07 51.74
@@
4.2 Experimental details
In CDGZSL, there are both seen and unseen domains, as well as seen and unseen classes. For Office-Home, the allocation of seen and unseen domains is as follows: Art, Clipart, and Product are each designated as the unseen domain in turn. When a domain is designated as the unseen domain, the other three domains serve as the seen domains. Because we employ Resnet-50 pretrained on Imagenet(Deng et al., 2009) as the initial parameters for the data encoder $E_{d}$, only the Real-world domain is not used as the unseen domain (since it is more similar to Imagenet) among these four domains. Additionally, the division of seen and unseen classes is as follows: we select 5 classes that do not appear in Imagenet as unseen classes, 5 classes as validation, and the remaining 55 classes as seen classes. For Mini-DomainNet, similarly, aside from the Real domain, the other three domains each take a turn as the unseen domain once. Also, we select 13 classes that do not appear in Imagenet as unseen classes, 13 classes as validation, and the remaining classes as seen classes.
For the various networks used in the experiments, the specific details are as follows: The data encoder $E_{d}$ is the encoder of Resnet50, with initial parameters trained on Imagenet. The domain classifier $C_{d}$, seen class classifier $C_{s}$, and the final CDGZSL classifier $C_{c}$ are all single-layer fully connected layers followed by Softmax. The semantic encoder $E_{s}$ is divided into two cases: When the semantics are text generated by LLM, $E_{s}$ is sentence-Bert (Reimers and Gurevych, 2019), with initial parameters using all-MiniLM-L6-v2, and only the last two layers are fine-tuned during the training. When the semantics are manually annotated attributes, $E_{s}$ is a two-layer fully connected layer with an output dimension of 15. The generator and discriminator of WGAN-GP are also both two-layer fully connected layers. The activation function of all the above-mentioned fully connected layers is ReLU. For the optimizers and various hyperparameters used in training, the specific details are as follows: $\lambda=1$, $T=5$, $\alpha=0.1$, $\beta=1$, the optimizer is Adam, and the learning rate is $2\times 10^{-4}$.